Let f (x) = x^2 + 4x - 7
(a) Find the coordinates of the vertex.
(x, y) = ___

(b) The [Either max/min] value of f is
f ( __ ) = __

(c) Find the domain and range of f

Answer by stanbon(75887)

Let f (x) = x^2 + 4x - 7
(a) Find the coordinates of the vertex.
(x, y) =
x = -b/(2a) = -4/(2*1) = -2
y = (-2)^2 + 4*-2 -7 = -11
----
(x,y) = (-2,-11)
-----------------------
(b) The [Either max/min] value of f is
f (-2)) = -11
-----------------------
(c) Find the domain and range of f
Domain:: All Real Numbers
Range:: y >= -11
--------------------
Cheers,
Stan H.